# MemANNS: Enhancing Billion-Scale ANNS Efficiency with Practical PIM Hardware

### v1

@@ 6. Related Work
Software-based Methods. Software-based methods primarily enhance the performance of ANNS through algorithm-level optimizations. For instance, VDTuner (Yang et al., 2024) utilizes a learning-based parameter tuning method to more quickly find the optimal parameter configuration for ANNS databases. Works such as  (André et al., 2016; LeJeune et al., 2019; Gao and Long, 2023) reduce memory access frequency by establishing a distance or using pruning methods. In contrast, our proposed method accelerates the ANNS algorithm based on new hardware, complementing the aforementioned software-based methods.
Hardware-based Methods. Hardware-based methods primarily accelerate ANNS by utilizing existing hardware or designing new hardware. GPUs are the most commonly used hardware to accelerate ANNS. For example, works such as (Zhao et al., 2020; Khan et al., 2024; Groh et al., 2022) utilize GPU-based graph methods to accelerate ANNS, but due to the limited and expensive memory capacity in GPUs, it is challenging to scale to large datasets. Juno (Liu et al., 2024) is based on a compression-based method that uses GPU’s RT-cores to accelerate the IVFPQ method. Compared to graph-based methods, compression-based methods can significantly reduce storage costs, making them better suited to the current situation of rapidly growing data scale. However, compared to MemANNS, these GPU-based methods have a lower energy cost-effectiveness.
In addition to GPUs, many other hardware options have also been used to accelerate ANNS. For example, works such as (Abdelhadi et al., 2019; Jiang et al., 2023) utilize FPGA to accelerate ANNS. Although this achieves high throughput on million-scale datasets, scaling to billion-scale datasets is limited by the on-chip memory of FPGAs. Conversely, MemANNS is well-suited for billion-scale datasets and offers easy scalability. The works such as (Kim et al., 2023; Tian et al., 2024) utilize SmartSSDs to accelerate ANNS. While this offers a large memory capacity, it has limited bandwidth and can exacerbate the communication bottleneck of PCIe when ANNS is used alongside other applications like recommendation systems and language models (LLMs). In contrast, PIM provides significantly larger and scalable bandwidth, which can communicate with the host via the memory bus without increasing the PCIe communication burden. In addition, the work (Jang et al., 2023) utilizes Compute Express Link(CXL) to accelerate ANNS by incorporating additional memory space and controllers in the main memory. However, this method can significantly increase the cost of system maintenance. In contrast, MemANNS is a cost-effective and easily scalable solution that offers near-linear performance scalability for large-scale datasets.
In addition to existing hardware, many works design new hardware specifically tailored to the characteristics of ANNS for acceleration. For example, (Lee et al., 2022) proposed a specialized architecture, which combines the benefits of a specialized dataflow pipeline and efficient data reuse to accelerate ANNS. The work (Wang et al., 2024) proposed a near-data processing (NDP) architecture based on SmartSSD to accelerate the graph-traversal-based ANNS task. However, the aforementioned works are still at the simulation or prototype stage, and are far from practical application.
@@ 7. Conclusion
In this work, we addressed the significant performance bottlenecks faced by CPU and GPU-based Approximate Nearest Neighbor Search (ANNS) solutions at billion-scale datasets, where CPU-based solutions are bounded by limited memory bandwidth and GPU-based solutions encounter memory capacity and resource utilization issues. We present MemANNS, a novel framework leveraging UPMEM’s Processing-in-Memory (PIM) architecture to overcome the memory bottleneck in billion-scale ANNS algorithms. MemANNS effectively addresses memory bottlenecks by employing architecture-aware data placement, efficient resource management, and a novel encoding approach for the IVFPQ algorithm. Our evaluation demonstrates that MemANNS significantly improves performance, achieving a remarkable 4.3x increase in QPS compared to CPU-based implementations of Faiss, while also matching the performance of GPU-based Faiss systems. Furthermore, MemANNS shows a commendable 2.3x improvement in QPS per Watt compared to GPU solutions, highlighting its superior cost-effectiveness and potential for large-scale applications such as large model serving.

### v2

@@ 6. Related Work
Software-based Methods. Software-based methods primarily enhance the performance of ANNS through algorithm-level optimizations. For instance, VDTuner (Yang et al., 2024) utilizes a learning-based parameter tuning method to more quickly find the optimal parameter configuration for ANNS databases. Works such as  (André et al., 2016; LeJeune et al., 2019; Gao and Long, 2023) reduce memory access frequency by establishing a distance bound or using pruning methods. In contrast, our proposed method accelerates the ANNS algorithm based on new hardware, complementing the aforementioned software-based methods.
Hardware-based Methods. Hardware-based methods primarily accelerate ANNS by utilizing existing hardware or designing new hardware. GPUs are the most commonly used hardware to accelerate ANNS. For example, works such as (Zhao et al., 2020; Khan et al., 2024; Groh et al., 2022) employ GPU-optimized graph algorithms for ANNS but face scalability challenges due to limited GPU memory. Juno (Liu et al., 2024) addresses this using GPU ray-tracing cores (RT-cores) to accelerate the IVFPQ compression-based method, reducing storage costs compared to graph-based methods. However, GPU solutions generally exhibit lower energy efficiency than UpANNS.
In addition to GPUs, many other hardware options have also been used to accelerate ANNS. For example, works such as (Abdelhadi et al., 2019; Jiang et al., 2023a; Yuan and Jin, 2025) utilize FPGA to achieve high throughput on million-scale datasets but struggle with billion-scale data due to constrained on-chip memory. SmartSSD-based systems (Kim et al., 2023; Tian et al., 2024) offer ample memory but risk PCIe bandwidth bottlenecks when coexisting with applications like LLMs. Compute Express Link (CXL) (Jang et al., 2023) accelerates ANNS by expanding memory via additional controllers, yet raises system maintenance costs. In contrast, UpANNS is a cost-effective and easily scalable solution that offers near-linear performance scalability for large-scale datasets.
In addition to existing hardware, many works design custom hardware specifically tailored to the characteristics of ANNS for acceleration. For example, (Lee et al., 2022a) proposed a specialized architecture, which combines the benefits of a specialized dataflow pipeline and efficient data reuse to accelerate ANNS. The work (Wang et al., 2024) proposed a near-data processing (NDP) architecture based on SmartSSD to accelerate the graph-traversal-based ANNS task. However, the aforementioned works are still at the simulation or prototype stage, and are far from practical application.
@@ 7. Conclusion
In this work, we addressed the significant performance bottlenecks faced by CPU and GPU-based Approximate Nearest Neighbor Search (ANNS) solutions at billion-scale datasets, where CPU-based solutions are bounded by limited memory bandwidth and GPU-based solutions encounter memory capacity and resource utilization issues. We present UpANNS, a novel framework leveraging UPMEM’s Processing-in-Memory (PIM) architecture to overcome the memory bottleneck in billion-scale ANNS algorithms. UpANNS effectively addresses memory bottlenecks by employing architecture-aware data placement, efficient resource management, and a novel encoding approach for the IVFPQ algorithm. Our extensive evaluation demonstrates that UpANNS significantly improves performance, achieving a remarkable 4.3x increase in QPS compared to CPU-based implementations of Faiss, while also matching the performance of GPU-based Faiss systems. Furthermore, UpANNS shows a commendable 2.3x improvement in QPS per Watt compared to GPU solutions, highlighting its superior cost-effectiveness and potential for large-scale applications such as large model serving.
Although our current implementation is tuned for IVFPQ, the core techniques, namely workload distribution, resource management, and top-k pruning, are transferable. Future work will generalize UpANNS to broader ANNS algorithms, and exploit next-generation PIM hardware with higher frequency and bandwidth to further improve competitiveness against high-end accelerators.